OpenStudy (anonymous):

What is the probability of getting either an ace or a five when drawing a single card from a deck of 52 cards? Formula P(a or b)= p(a) + P(b)- p(A and B)

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B) I don't think that this is right??? HELP! P(A or B) = 4/52 + 4/52 - 8/52 the answer would be 0

OpenStudy (anonymous):

I am not good with probabilities...but I am going to give it a try... probability of getting an ace is 4/52 (because there is 4 aces) which reduces to 1/13 probability of getting a 5 is 4/52(because there is 4 fives) which reduces to 1/13 I think the probability that both can occur is 1/13 x 1/13 = 2/169 you might need to get a second opinion on this

OpenStudy (kropot72):

The events 'draw an ace' and 'draw a five' are mutually exclusive (you cannot draw both an ace and a five when drawing a single card). Therefore the correct formula to use is \[P(A \cup B)=P(A)+P(B)\] This formula gives the probability of A or B occurring for mutually exclusive events. The formula given by @kimbradunn gives the probability of event A or B (or both).

OpenStudy (anonymous):

oh...so you add them....I multiplied them....I told you I wasn't good at this.

OpenStudy (kropot72):

np